You receive 14 tv's. Two of the tv's are defective. If two tv's are randomly selected, compute the probability that both tv's work. What is the probability at least one does not work?
-The probability that both tv's work is -
-The probability that at least one does not work is -

Answer by stanbon(75887)

You receive 14 tv's. Two of the tv's are defective. If two tv's are randomly selected, compute the probability that both tv's work. What is the probability at least one does not work?
P(a selected tv is defective) = 2/14 = 1/7
P(a selected tv is not defective) = 6/7
----------------------------------------------
-The probability that both tv's work is -
P(2 work) = 2C2/14C2 = 1/91
-------------------------------------
-The probability that at least one does not work is -
P(at least one is bad) = 1 -P(none is bad)
= 1 - 12C2/14C2
= 1 - 66/91
= 25/91
================
Cheers,
Stan H.